OpenStudy (anonymous):

Let triangle A B C is similar to triangle D E F. In triangle a b c, side a b equals five and side b c equals ten. side d e is unknown, and side e f equals eight. What is the measure of side d e ? 4 8 10 16

Directrix (directrix):

Lengths of corresponding sides of similar triangles are in proportion. AB/DE = BC/EF 5/DE = 10/8 DE = 4

OpenStudy (anonymous):

Thank you but I already found the answer... =)

Directrix (directrix):

@SavannahSaturn --> I enjoy working on the problems and rarely notice if they are open or closed. Regardless, I hope we got the same answer. :)

OpenStudy (anonymous):

Yes we did and I actually got it right. Im kinda the same way. I really enjoy geometry so I have no problem doing it myself sometime. =)